One number is 2 greater than another. The product of the numbers is 63. Find the numbers.
Answer by ikleyn(52835)

.
One number is 2 greater than another. The product of the numbers is 63. Find the numbers.
~~~~~~~~~~~~~~~~~~~~~~~~~~

Answer. a) 7 and 9; b) -9 and -7.

Solution

x*(x+2) = 63,

x^2 + 2x -63 = 0,

Factor left side

(x+9)*(x-7) = 0.

The roots are 

  a)  x = 7.  Then x+2 = 9.

      or

  b)  x = -9.  Then  x+2 = -7.

Solved.